Walton Manufacturing Company was started on January 1, 2018, when it acquired $90,000 cash by issuing common stock. Walton immediately purchased office furniture and manufacturing equipment costing $7,000 and $32,200, respectively. The office furniture had an eight-year useful life and a zero salvage value. The manufacturing equipment had a $3,000 salvage value and an expected useful life of four years. The company paid $11,300 for salaries of administrative personnel and $15,900 for wages to production personnel. Finally, the company paid $11,300 for raw materials that were used to make inventory. All inventory was started and completed during the year. Walton completed production on 4,600 units of product and sold 3,680 units at a price of $14 each in 2018. (Assume that all transactions are cash transactions and that product costs are computed in accordance with GAAP.)

Required

  1. Determine the total product cost and the average cost per unit of the inventory produced in 2018. (Round "Average cost per unit" to 2 decimal places.)

  2. Determine the amount of cost of goods sold that would appear on the 2018 income statement. (Do not round intermediate calculations.)

  3. Determine the amount of the ending inventory balance that would appear on the December 31, 2018, balance sheet. (Do not round intermediate calculations.)

  4. Determine the amount of net income that would appear on the 2018 income statement. (Round your answer to the nearest dollar amount.)

  5. Determine the amount of retained earnings that would appear on the December 31, 2018, balance sheet. (Round your answer to the nearest dollar amount.)

  6. Determine the amount of total assets that would appear on the December 31, 2018, balance sheet. (Round your answer to the nearest dollar amount.)

I need to find the Total Product Cost, average cost per unit, cost of sold goods,ending inventory, net income,retained earnings,total assests... can you help me get these and understand how
Answer #1

Ans:

Total product cost and average cost:

Direct Materials= 11,300

Direct labor= 15,900

Manufacturing overhead (32,200 -3,000)/4 = 7,300

Total product cost= 11,300+15,900+7,300

= $ 34,500

Units= 4,600

Average cost= 7.50 per unit

B). Amount of cost of goods sold that would appear on income statement:

Number of units sold= 3,680

Average cost= 7.50

Cost of goods sold= 3,680*7.50

= 27,600

C).Amount of ending inventory balance that would appear on dec 31,2018=

Number of units= 4,600

Number of units sold= 3,680

Ending inventory units= 920

Average cost per unit= 7.50

Ending inventory Balance= 920*$7.50

= $6,900

D). Amount of net income that would appear on income statement

Sales= (3,680*14) : 51,520

Less: Cost of goods sold= 27,600

Administrative salaries= 11,300

Depreciation on office furniture= (7,000/8) = 875

Net income=11,745

E). Amount of retained earnings=

Beginning retained earnings= 0

Add: net income= 11,745

Less: dividends= 0

Ending retained earnings= 11,745

F). Amount of total assets that would appear on Dec 31,2018=

Cash= (90,000-7,000-32,200-11,300-15,900-11,300+51,520)= 63,820

Ending inventory= 6,900

office equipment= 7,000-3,000=4,000

Manufacturing equipment = (32,200-7,300)= 24,900

Total Assets= 63,820+6,900+4,000+24,900

= 99,620